### v1

@@ 3.1 Automatic Instance Segmentation
We observe that PathoSAM with AIS is the best overall model, despite our reporting favoring methods with multiple versions. On the ”in-domain“ datasets, PathoSAM is either the best or on par with the best method. For the ”out-of-domain“ datasets it shows remarkable generalization, outperforming other methods for NuInsSeg (H&E-stained images) and IHC-TMA (IHC-stained tissue microarrays). CryoNuSeg (cryo-sectioned H&E-stained images) is the only dataset where it underperforms, see Fig. 5 for details. We include the comparison to PathoSAM (PanNuke), to investigate the advantage of training on multiple datasets. The effectiveness of this approach is validated, as the generalist model is on par or better for all datasets except IHC-TMA. We also see that AIS is supperior to AMG in most cases. Hence this approach is clearly preferable, also due to its better efficiency [Archit et al.(2023)Archit, Nair, Khalid, Hilt, Rajashekar, Freitag, Gupta, Dengel, Ahmed, and Pape]. Finally, the results in 7 show that differences between ViT-B, ViT-L and ViT-H are small, with ViT-L reaching the highest score.
@@ 3.2 Interactive Instance Segmentation
We compare PathoSAM (generalist, ViT-B), with other SAM variants for interactive segmentation. Specifically, we compare it with the original SAM [Kirillov et al.(2023)Kirillov, Mintun, Ravi, Mao, Rolland, Gustafson, Xiao, Whitehead, Berg, Lo, Dollar, and Girshick] and with $\mu$SAM [Archit et al.(2023)Archit, Nair, Khalid, Hilt, Rajashekar, Freitag, Gupta, Dengel, Ahmed, and Pape], which was finetuned on a large microscopy dataset. We use the same 12 datasets as before. The results for 8 selected datasets are reported in Fig. 2 b) and the overall results in Fig. 1 c). Detailed results are given in App. D.2. We use the evaluation procedure introduced by $\mu$SAM, see also Sec. 2.1, reporting the mean segmentation accuracy for segmentation with a point prompt, a box prompt, iterative correction after an initial point, and iterative correction after an initial box.
@@ 3.3 Other Segmentation Tasks
We also evaluate PathoSAM on three other segmentation tasks. First, we evaluate segmentation of different structures, lymphocytes in IHC-stained tissue images [Alemi Koohbanani et al.(2020)Alemi Koohbanani, Jahanifar, Zamani Tajadin, and Rajpoot] and glands in H&E-stained colon tissue images [Sirinukunwattana et al.(2016)Sirinukunwattana, Pluim, Chen, Qi, Heng, Guo, Wang, Matuszewski, Bruni, Sanchez, Böhm, Ronneberger, Cheikh, Racoceanu, Kainz, Pfeiffer, Urschler, Snead, and Rajpoot]. Here, we compare the automatic and interactive segmentation with SAM, $\mu$SAM, and PathoSAM. We also finetune a new model, PathoSAM (specialist), on the training splits of the respective datasets. Quantitative and qualitative results are shown in Fig. 3 a) and b). For lymphocytes, SAM, $\mu$SAM and PathoSAM (generalist) cannot automatically segment the data well, but provide good interactive segmentation. Conversely, the specialist model yields better automatic segmentation, but is worse for interactive segmentation. The reason for the difference in automatic segmentation is clear: the lymphocytes only make up a fraction of the nuclei in the images. In automatic segmentation, the non-specific models segment all nuclei, resulting in a bad score. The specialist model distinguishes lymphocytes from other nuclei, thus providing a better result. It is unclear why it performs worse in interactive segmentation. Overall, this task may be better formulated as a semantic instance segmentation problem, see also next paragraph, to avoid penalizing the segmentation of all nuclei. For gland segmentation, the specialist outperforms all other SAM variants. The glands have a different appearance and size compared to nuclei, see the images on the right, explaining this observation. In both cases, the specialist was trained based on the default SAM, which provided the best results among the non-specific models. Overall, these experiments show that neither of the available SAM variants are sufficient to address general instance segmentation tasks beyond nuclei. Finetuning for a specific task using our recipe can however be used to obtain a good tailored model.
@@
We also evaluate PathoSAM for semantic nucleus segmentation on PanNuke [Gamper et al.(2020)Gamper, Koohbanani, Benes, Graham, Jahanifar, Khurram, Azam, Hewitt, and Rajpoot]. For this application, we train an additional decoder based on the PathoSAM generalist, see Sec. 2.4 for details. The comparison to other methods trained for this task on PanNuke are shown in Fig. 3 c). Here, we observe that PathoSAM is the second best model, after CellViT, the clear state-of-the-art. HoverNet and HoverNext are on par and better than BioMedParse. Additional results for semantic segmentation are given in 8 and App. D.1, reporting the individual class scores and, for PathoSAM, the result for different training strategies. Here, the main observation is that only CellViT correctly segments the “dead cells” minority class (<2 % of the nuclei). Furthermore, finetuning the encoder and training the new decoder from scratch yields the best results for PathoSAM.
We also conduct a user study to assess the practical use of PathoSAM. Here, we check the integration of PathoSAM with two user-friendly tools: QuPath [Bankhead et al.(2017)Bankhead, Loughrey, Fernández, Dombrowski, McArt, Dunne, McQuaid, Gray, Murray, Coleman, James, Salto-Tellez, and Hamilton], using the SAM-API plugin [Sugawara(2023)], and $\mu$SAM [Archit et al.(2023)Archit, Nair, Khalid, Hilt, Rajashekar, Freitag, Gupta, Dengel, Ahmed, and Pape]. We use the tools to annotate nuclei in [Janowczyk and Madabhushi(2016)], which contains nuclei in H&E stained breast tissue images. The qualitative results, comparing SAM and PathoSAM, are shown in Fig. 11 and 12 respectively. We see clear improvements due to PathoSAM in both cases, as nuclei are much better captured from single point or box prompts with it. Beyond tool integration, we also offer scripts to apply our model to whole-slide images to enable automatic segmentation in practice.
@@ 4 Discussion
Our model, PathoSAM, is the new state-of-the-art for interactive and automatic instance segmentation of nuclei in histopathology. To our knowledge, it is the first foundation model for this task, as it generalizes to diverse settings with a single model. Hence, we believe that it will have great practical impact by speeding up many nucleus analysis tasks. To support practical use, we provide integration with user-friendly tools and scripts for batch processing of whole-slide images.
We have also evaluated PathoSAM for segmentation of lymphocytes and glands, where we found that PathoSAM does not provide benefits over other SAM variants, but have shown that finetuning with our recipe can yield a good tailored model. Tool integration (see above) of our model enables fast annotation in such cases. For semantic segmentation we found that further finetuing PathoSAM yields good results, but that it is not yet competitive with CellViT, the state-of-the-art. Notably, we have only evaluated this task on a single dataset, so it is unclear how any of the methods generalize.. A more diverse dataset, from BioMedParse [Zhao et al.(2024)Zhao, Gu, Yang, Usuyama, Lee, Kiblawi, Naumann, Gao, Crabtree, Abel, Moung-Wen, Piening, Bifulco, Wei, Poon, and Wang], exists. This could enable training and evaluation of a VFM that also supports semantic segmentation. PathoSAM could then also be improved by implementing a joint training procedure with instance segmentation, similar to CellViT.
@@
The work of Titus Griebel was supported by the Göttingen Promotionskolleg für Medizinstudierende, funded by Else Kröner-Fresenius-Stiftung. The work of Anwai Archit was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - PA 4341/2-1. This work was also supported by the Google Research Scholarship “Vision Foundation Models for Bioimage Segmentation”. We gratefully acknowledge the computing time granted by the Resource Allocation Board and provided on the supercomputer Emmy at NHR@Göttingen as part of the NHR infrastructure, under the project nim00007. We would like to thank Sebastian von Haaren for suggestions on improving data visualizations.
@@ Appendix A Evaluation Metric
We use the mean segmentation accuracy, following [Caicedo et al.(2019)Caicedo, Goodman, Karhohs, Cimini, Ackerman, Haghighi, Heng, Becker, Doan, McQuin, Rohban, Singh, and Carpenter], to evaluate instance segmentation results. It is based on true positives ($TP$), false negatives ($FN$), and false positive ($FP$), which are derived from the intersection over union (IoU) of predicted and true objects. Specifically, $TP(t)$ is defined as the number of matches between predicted and true objects with an IoU above the threshold $t$, $FP(t)$ correspond to the number of predicted objects minus $TP(t)$, and $FN(t)$ to the number of true objects minus $TP(t)$. The mean segmentation accuracy is computed over multiple thresholds:
@@
We use the dice coefficient to evaluate the semantic segmentation results. It is defined as
@@
Note that prior work regularly reports the panoptic quality (PQ) as metric for semantic instance segmentation on PanNuke, the dataset we use to evaluate the semantic segmentation. However, we reviewed the referenced code at https://github.com/TissueImageAnalytics/PanNuke-metrics/blob/master/utils.py and found that it is not a proper implementation of the PQ, and that it seems to only evaluate the semantic segmentation. Hence, we decided to use the Dice coefficient, which is a more established metric for semantic segmentation. We will investigate this further during the revision, and if possible will report a true PQ metric.
@@ Appendix B Dataset Description
Table 1: Description of the different datasets used in our study for training generalist models, evaluating and comparing them to other methods, and evaluating downstream tasks.
@@ Appendix C Architecture and Training Details
We train the generalist model with the objective from $\mu$SAM [Archit et al.(2023)Archit, Nair, Khalid, Hilt, Rajashekar, Freitag, Gupta, Dengel, Ahmed, and Pape] that combines training of interactive and automatic segmentation (see Sec. 2.1 and Sec. 2.2). The generalist models are trained for 250,000 iterations, the PanNuke model and the specialist models for downstream tasks are trained for 100,00 iterations. We use 7 correction iterations in the training objective and a probability for using a mask prompt of 50% per iteration. This choice enables interactive both with and without addition of the mask prompt. We use a batch size of 2 and train with 40 (ViT-B), 30 (ViT-L), or 25 (ViT-H) objects per image in the batch as hyperparameters for the training objective. We use the AdamW optimizer with an initial learning rate of 1e-5. We use a learning rate scheduler that scales it by a factor of 0.9 when the validation metric plateuas for 10 epochs. All models are trained on a single NVIDIA A100 GPU with 80GB VRAM.
@@ D.1 Automatic Segmentation
We provide comprehensive overview of all results for automatic instance segmentation. 6 displays results for all 8 datasets discussed in the main paper, and additional 5 datasets, which include CPM15 [Vu et al.(2019)Vu, Graham, Kurc, To, Shaban, Qaiser, Koohbanani, Khurram, Kalpathy-Cramer, Zhao, Gupta, Kwak, Rajpoot, Saltz, and Farahani], CPM17 [Vu et al.(2019)Vu, Graham, Kurc, To, Shaban, Qaiser, Koohbanani, Khurram, Kalpathy-Cramer, Zhao, Gupta, Kwak, Rajpoot, Saltz, and Farahani], CoNSeP [Graham et al.(2019)Graham, Vu, Raza, Azam, Tsang, Kwak, and Rajpoot], LyNSeC (IHC) [Naji et al.(2024)Naji, Sancere, Simon, Büttner, Eich, Lohneis, and Bożek] and TNBC [Naylor et al.(2019)Naylor, Laé, Reyal, and Walter]. Further, 7 includes detailed quantitative results for all the datasets. We provide results for all combination of models for the benchmarked methods, including all trained PathoSAM generalist models. We observe that PathoSAM is consistently either the best or among the best models for the datasets, including for out-of-domain datasets that contain different stainings (e.g. IHC staining) or imaging conditions (e.g. CryoNuSeg). For CryoNuSeg, which contains images of tissues that were cryo-sectioned, the automatic segmentation quality of the generalist model suffers from this domain shift. They also have low scores for the automatic segmentation of lymphocytes in NuClick, which is due to the fact that other nuclei are segmented as well. This dataset is used for further investigation in Sec. 3.3 to understand if training a tailored specialist model leads to an improved automatic lymphocytes segmentation.

### v2

@@ 3.1 Automatic Instance Segmentation
We observe that PathoSAM with AIS is the best overall model, despite our reporting favoring methods with multiple versions. On the ”in-domain“ datasets, PathoSAM is either the best or on par with the best method. For the ”out-of-domain“ datasets it shows remarkable generalization, outperforming other methods for NuInsSeg (H&E-stained images) and IHC-TMA (IHC-stained tissue microarrays). CryoNuSeg (cryo-sectioned H&E-stained images) is the only dataset where it underperforms, see Fig. 6 for details. We include the comparison to PathoSAM (PanNuke), to investigate the advantage of training on multiple datasets. The effectiveness of this approach is validated, as the generalist model is on par or better for all datasets except IHC-TMA. We also see that AIS is superior to AMG in most cases. Hence this approach is clearly preferable, also due to its better efficiency [Archit et al.(2025a)Archit, Freckmann, Nair, Khalid, Hilt, Rajashekar, Freitag, Teuber, Buckley, von Haaren, Gupta, Dengel, Ahmed, and Pape]. The results in 8 show that differences between ViT-B, ViT-L and ViT-H are small, with ViT-L reaching the highest score. Furthermore, we investigate the segmentation quality across overlap thresholds in Fig. 9 and the influence of post-processing parameters on the precision-recall trade-off in Fig. 10. We also study the segmentation results for neutrophils, which have complex nuclear shapes, in Fig. 11.
@@ 3.2 Interactive Instance Segmentation
We compare PathoSAM (generalist, ViT-B), with other SAM variants for interactive segmentation. Specifically, we compare it with the original SAM [Kirillov et al.(2023)Kirillov, Mintun, Ravi, Mao, Rolland, Gustafson, Xiao, Whitehead, Berg, Lo, Dollar, and Girshick] and with $\mu$SAM [Archit et al.(2025a)Archit, Freckmann, Nair, Khalid, Hilt, Rajashekar, Freitag, Teuber, Buckley, von Haaren, Gupta, Dengel, Ahmed, and Pape], which was finetuned on a large microscopy dataset. We use the same 12 datasets as before. The results for 8 selected datasets are reported in Fig. 2 b) and the overall results in Fig. 1 c). Detailed results are given in App. D.2. We use the evaluation procedure introduced by $\mu$SAM, see also Sec. 2.1, reporting the mean segmentation accuracy for segmentation with a point prompt, a box prompt, iterative correction after an initial point, and iterative correction after an initial box.
@@ 3.3 Other Segmentation Tasks
We also evaluate PathoSAM on three other segmentation tasks. First, we evaluate segmentation of different structures, lymphocytes in IHC-stained tissue images [Alemi Koohbanani et al.(2020)Alemi Koohbanani, Jahanifar, Zamani Tajadin, and Rajpoot] and glands in H&E-stained colon tissue images [Sirinukunwattana et al.(2017)Sirinukunwattana, Pluim, Chen, Qi, Heng, Guo, Wang, Matuszewski, Bruni, Sanchez, Böhm, Ronneberger, Cheikh, Racoceanu, Kainz, Pfeiffer, Urschler, Snead, and Rajpoot]. Here, we compare the automatic and interactive segmentation with SAM, $\mu$SAM, and PathoSAM. We also finetune a new model, PathoSAM (specialist), on the training splits of the respective datasets. Quantitative and qualitative results are shown in Fig. 3 a) and b). For lymphocytes, SAM, $\mu$SAM and PathoSAM (generalist) cannot automatically segment the data well, but provide good interactive segmentation. Conversely, the specialist model yields better automatic segmentation, but is worse for interactive segmentation. The reason for the difference in automatic segmentation is clear: the lymphocytes only make up a fraction of the nuclei in the images. In automatic segmentation, the non-specific models segment all nuclei, resulting in a bad score. The specialist model distinguishes lymphocytes from other nuclei, thus providing a better result. It is unclear why it performs worse in interactive segmentation. Overall, this task may be better formulated as a semantic instance segmentation problem, see also next paragraph, to avoid penalizing the segmentation of all nuclei. For gland segmentation, the specialist outperforms all other SAM variants. The glands have a different appearance and size compared to nuclei, see the images on the right, explaining this observation. In both cases, the specialist was trained based on the default SAM, which provided the best results among the non-specific models. Another architecture based on SAM has been proposed for gland segmentation [Zhang et al.(2024)Zhang, Li, Xue, Wang, and Li]. However, this approach addresses semantic segmentation, not instance segmentation, so we do not compare to it here. Overall, these experiments show that neither of the available SAM variants are sufficient to address general instance segmentation tasks beyond nuclei. Finetuning for a specific task using our recipe can however be used to obtain a good tailored model.
@@
We also evaluate PathoSAM for semantic nucleus segmentation on PanNuke [Gamper et al.(2019)Gamper, Koohbanani, Benes, Khuram, and Rajpoot]. For this application, we train an additional decoder based on the PathoSAM generalist, see Sec. 2.4 for details. The comparison to other methods trained for this task on PanNuke are shown in Fig. 3 c). Here, we observe that PathoSAM is the second best model, after CellViT, the clear state-of-the-art. HoverNet and HoverNext are on par and better than BioMedParse. Additional results for semantic segmentation are given in 12 and App. D.1, reporting the individual class scores and, for PathoSAM, the result for different training strategies. Here, the main observation is that only CellViT correctly segments the “dead cells” minority class (<2 % of the nuclei). Furthermore, finetuning the encoder and training the new decoder from scratch yields the best results for PathoSAM.
@@
3.4 Practical Use of PathoSAM
We evaluate the practical utility of PathoSAM by conducting a qualitative user study for interactive data annotation. Here, we check integration with two user-friendly tools: QuPath [Bankhead et al.(2017)Bankhead, Loughrey, Fernández, Dombrowski, McArt, Dunne, McQuaid, Gray, Murray, Coleman, James, Salto-Tellez, and Hamilton], using the SAM-API plugin [Sugawara(2023)], and $\mu$SAM [Archit et al.(2025a)Archit, Freckmann, Nair, Khalid, Hilt, Rajashekar, Freitag, Teuber, Buckley, von Haaren, Gupta, Dengel, Ahmed, and Pape]. We use the tools to annotate nuclei in [Janowczyk and Madabhushi(2016)], which contains nuclei in H&E stained breast tissue images. The qualitative results, comparing SAM and PathoSAM, are shown in Fig. 15 and 16. We see clear improvements with PathoSAM in both tools: it captures nuclei better from single point or box prompts.
We also implement tile-and-stitch based prediction to apply PathoSAM to whole-slide images (WSI) and enable automatic segmentation in practice. By default, we use a tile shape of 512 $\times$ 512 pixels, which corresponds to the training patch shape of PathoSAM, with an overlap of 64 pixels on each side. See Fig. 4 for an example and for an overview of the resources needed for WSI inference. Segmentation of a WSI runs in less than 1 hour on a GPU, requiring ca. 5 GB of VRAM per tile. It takes much longer on a CPU (ca. 37 hrs). Our implementation on the GPU has been optimized. The encoder and decoder are applied with a batch size of 1 on CPU resource, otherwise it is automatically selected depending on available VRAM. Increasing the batch size to process multiple tiles in parallel speeds up inference, especially on GPUs with large VRAM.
@@ 4 Discussion
Our model, PathoSAM, is the new state-of-the-art for interactive and automatic instance segmentation of nuclei in histopathology. To our knowledge, it is the first foundation model for this task, as it generalizes to diverse settings with a single model. We believe that it will have great practical impact by speeding up many nucleus analysis tasks. To support this use, we provide integration with user-friendly tools and scripts for processing WSIs.
Limitations:
PathoSAM does not improve segmentation for other tasks in histopathology, e.g. gland or lymphocyte segmentation (see Sec.  3.3). It is thus not yet a general purpose foundation model for histopathology segmentation, but only for nucleus segmentation. However, we show that our finetuning methodology can be used to obtain improved models for other tasks as well; our tool integration provides fast data annotation for such cases. For semantic segmentation we found that PathoSAM yields good results, but that it is not yet competitive with CellViT, the state-of-the-art. We believe that this is due to the sophisticated procedure CellViT uses to oversample patches containing minority class annotations. In contrast, we train without any over-sampling or loss balancing. Notably, we have only evaluated this task on a single dataset, so it is unclear how any of the methods for semantic segmentation generalize.
Future Work:
We plan to improve PathoSAM for semantic segmentation by (i) implementing a joint training procedure for instance and semantic segmentation, (ii) over-sampling minority classes (see prev. paragraph), and (iii) assembling a large dataset for semantic training, including annotations from [Schuiveling et al.(2025)Schuiveling, Liu, Eek, Breimer, Suijkerbuijk, Blokx, and Veta] and [Graham et al.(2019)Graham, Vu, Raza, Azam, Tsang, Kwak, and Rajpoot]. This would provide a versatile foundation model for both instance and semantic segmentation. Furthermore, we will explore training foundation models based on SAM2 [Ravi et al.(2025)Ravi, Gabeur, Hu, Hu, Ryali, Ma, Khedr, Rädle, Rolland, Gustafson, Mintun, Pan, Alwala, Carion, Wu, Girshick, Dollar, and Feichtenhofer]. Ultimately, we would like to provide a foundation model that addresses further segmentation tasks, such as gland or tissue level segmentation, within a single model. This would however require a multi-scale segmentation approach to exploit tissue-level features that are more prominent after “zooming out”.
@@
fig:wsi_seg Refer to caption
Figure 4: Segmentation on WSI. The tile shape is 512 $\times$ 512, with an overlap of 64 pixels. a) shows an example for automatic instance segmentation on data from https://openslide.cs.cmu.edu/download/openslide-testdata/ (32,914 $\times$ 46,000 pixels). b) gives resource demand and runtime for the different inference steps using GPU (tested with an A100 with 80GB VRAM) and CPU. We indicate the resources available for the CPU; whereas the entire GPU VRAM available is utilized using batch size of 30. The minimal GPU VRAM requirement is 5GB.
[…]
The work of Titus Griebel was supported by the Göttingen Promotionskolleg für Medizinstudierende, funded by Else Kröner-Fresenius-Stiftung. The work of Anwai Archit was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - PA 4341/2-1. Constantin Pape is supported by the German Research Foundation (Deutsche Forschungsgemeinschaft, DFG) under Germany’s Excellence Strategy - EXC 2067/1-390729940. This work is supported by the Ministry of Science and Culture of Lower Saxony through funds from the program zukunft.niedersachsen of the Volkswagen Foundation for the ’CAIMed – Lower Saxony Center for Artificial Intelligence and Causal Methods in Medicine’ project (grant no. ZN4257). It was also supported by the Google Research Scholarship “Vision Foundation Models for Bioimage Segmentation”. We gratefully acknowledge the computing time granted by the Resource Allocation Board and provided on the supercomputer Emmy at NHR@Göttingen as part of the NHR infrastructure, under the project nim00007. We would like to thank Sebastian von Haaren for suggestions on improving data visualizations.
@@ Appendix A Evaluation Metric
We use the mean segmentation accuracy, following [Caicedo et al.(2019)Caicedo, Goodman, Karhohs, Cimini, Ackerman, Haghighi, Heng, Becker, Doan, McQuin, Rohban, Singh, and Carpenter], to evaluate instance segmentation results. It is based on true positives ($TP$), false negatives ($FN$), and false positives ($FP$), which are derived from the intersection over union (IoU) of predicted and true objects. Specifically, $TP(t)$ is defined as the number of matches between predicted and true objects with an IoU above the threshold $t$, $FP(t)$ correspond to the number of predicted objects minus $TP(t)$, and $FN(t)$ to the number of true objects minus $TP(t)$. The mean segmentation accuracy is computed over multiple thresholds:
@@
We use the dice coefficient to evaluate the semantic segmentation results. It is defined as:
@@
Note that prior work regularly reports the panoptic quality (PQ) as metric for semantic instance segmentation on PanNuke, the dataset we use to evaluate the semantic segmentation. However, we reviewed the referenced code at https://github.com/TissueImageAnalytics/PanNuke-metrics/blob/master/utils.py and found that it is not a proper implementation of the PQ, and that it seems to only evaluate the semantic segmentation. Hence, we decided to use the Dice coefficient, which is a more established metric for semantic segmentation. We will investigate this further in future work, and if possible will introduce a true PQ metric.
@@ Appendix B Dataset Description
Table 1: Description of the different datasets used in our study for training generalist models (marked in italics), evaluating (marked in bold) and comparing them to other methods, and evaluating downstream tasks (marked in underline).
@@ Appendix C Architecture and Training Details
We train the generalist model with the objective from $\mu$SAM [Archit et al.(2025a)Archit, Freckmann, Nair, Khalid, Hilt, Rajashekar, Freitag, Teuber, Buckley, von Haaren, Gupta, Dengel, Ahmed, and Pape] that combines training of interactive and automatic segmentation (see Sec. 2.1 and Sec. 2.2). The generalist models are trained for 250,000 iterations, the PanNuke model and the specialist models for downstream tasks are trained for 100,00 iterations. We use 7 correction iterations in the training objective and a probability for using a mask prompt of 50% per iteration. This choice enables interactive both with and without addition of the mask prompt. We use a batch size of 2 and train with 40 (ViT-B), 30 (ViT-L), or 25 (ViT-H) objects per image in the batch as hyperparameters for the training objective. We use the AdamW optimizer with an initial learning rate of 1e-5. We use a learning rate scheduler that scales it by a factor of 0.9 when the validation metric plateuas for 10 epochs. All models are trained on a single NVIDIA A100 GPU with 80GB VRAM.
@@ D.1 Automatic Segmentation
We provide comprehensive overview of all results for automatic instance segmentation. 7 displays results for all 8 datasets discussed in the main paper, and additional 5 datasets, which include CPM15 [Vu et al.(2019)Vu, Graham, Kurc, To, Shaban, Qaiser, Koohbanani, Khurram, Kalpathy-Cramer, Zhao, Gupta, Kwak, Rajpoot, Saltz, and Farahani], CPM17 [Vu et al.(2019)Vu, Graham, Kurc, To, Shaban, Qaiser, Koohbanani, Khurram, Kalpathy-Cramer, Zhao, Gupta, Kwak, Rajpoot, Saltz, and Farahani], CoNSeP [Graham et al.(2019)Graham, Vu, Raza, Azam, Tsang, Kwak, and Rajpoot], LyNSeC (IHC) [Naji et al.(2024)Naji, Sancere, Simon, Büttner, Eich, Lohneis, and Bożek] and TNBC [Naylor et al.(2019)Naylor, Laé, Reyal, and Walter]. Further, 8 includes detailed quantitative results for all the datasets. We provide results for all combination of models for the benchmarked methods, including all trained PathoSAM generalist models. We observe that PathoSAM is consistently either the best or among the best models for the datasets, including for out-of-domain datasets that contain different stainings (e.g. IHC staining) or imaging conditions (e.g. CryoNuSeg). For CryoNuSeg, which contains images of tissues that were cryo-sectioned, the automatic segmentation quality of the generalist model suffers from this domain shift. They also have low scores for the automatic segmentation of lymphocytes in NuClick, which is due to the fact that other nuclei are segmented as well. This dataset is used for further investigation in Sec. 3.3 to understand if training a tailored specialist model leads to an improved automatic lymphocytes segmentation.
@@
[…]
fig:autoseg-f1-precision-recall Refer to caption
Figure 9: F1 Score, Recall and Precision for IoU thresholds from 0.5 to 0.95, for all datasets, for automatic instance segmentation with PathoSAM model (Generalist - ViT-L). Names of datasets in italics are in-domain, and bold are out-of-domain.
[…]
fig:heatmap-lynsec-he Refer to caption
Figure 10: Grid search results over (the most important) parameters of the watershed used to obtain the instance segmentation in AIS. These parameters are the thresholds applied to center and boundary distances predicted by the segmentation decoder in order to find seeds. Please refer to [Archit et al.(2025a)Archit, Freckmann, Nair, Khalid, Hilt, Rajashekar, Freitag, Teuber, Buckley, von Haaren, Gupta, Dengel, Ahmed, and Pape] for details on the instance segmentation methodology. We report the F1 Score, recall and precision for a grid-search over the two thresholds on 15 validation images of the LyNSeC (H&E) dataset, i.e. an OOD dataset. We see that precision is best for small (boundary) threshold values whereas recall is best for large thresholds. This is due to the fact that large threshold values lead to multiple seeds being derived from the distance maps, resulting in over-segmentation. Low threshold values have the opposite effect and result in under-segmentation. Hence, tuning the (boundary) threshold enables varying the degree of over-/under-segmentation.
[…]
fig:neutrophils-seg Refer to caption
Figure 11: Qualitative and quantitative results of PathoSAM on neutrophil granulocytes, and a further specialist model trained on a few images for neutrophil granulocyte segmentation on [Schuiveling et al.(2025)Schuiveling, Liu, Eek, Breimer, Suijkerbuijk, Blokx, and Veta]. Neutrophils have multi-lobed nuclei, which are more challenging to segment compared to “regular” nuclei. a) The scores are on par, with a slight favor for the specialist model. Note: we report scores specifically for neutrophil granulocyte instance segmentation, excluding other nucleus types from the evaluation. b) An example image with annotations for neutrophils. The zoom-ins show the segmentation results for the generalist model (middle) and the neutrophil specialist model (right). Arrows indicate neutrophils that are segmented better by the specialist compared to the generalist. These results indicate that both models can segment neutrophils reasonably well, and that the specialist model performs a bit better than our generalist. Importantly, this shows that our method can segment such complex shapes correctly, and that in practice the quality of results depends on the presence and quality of the respective annotations in the training data.